# GraphAny: A Foundation Model for Node Classification on Any Graph

### v2

@@ 3.2 Learning Inductive Attention over LinearGNNs
While there are various ways to parameterize this attention module, finding an inductive solution is non-trivial since neural networks can easily fit some information specific to the training graph. We notice a necessary condition for an inductive attention is that it should be robust to transformations on features and labels, such as permutation or masking on some dimensions (Figure 4). This requires our attention module to be permutation invariant and robust to dimension changes, which motivates our design of distance features and entropy normalization respectively.
Permutation-Invariant Attention with Distance Features. We would like to design an attention module that is permutation invariant [7] along the data dimension. [^1] It is important to distinguish between domain symmetry (in this case, permutation of the nodes of the graph) and data symmetry (permutation of the feature and label dimensions). Invariance to domain symmetry (node permutations) is provided by design in our LinearGNN. Consider a new graph generated by permuting feature and label dimensions of the training graph. We expect our attention output to be invariant to these permutations in order to generate the same prediction as for the training set.
@@
which is invariant to both feature- and label-permutation matrices ${\bm{P}}$ and ${\bm{Q}}$. Generally, any feature that is a linear combination of dot products between LinearGNN predictions is also permutation invariant, e.g. Euclidean distance or Jensen-Shannon divergence. For a set of LinearGNN predictions $\hat{{\bm{y}}}_{u}^{(1)},\hat{{\bm{y}}}_{u}^{(2)},\cdots,\hat{{\bm{y}}}_{u}^{(t)}$ on a single node $u$, we construct the following $t(t-1)$ permutation-invariant features to capture their squared difference:
@@
Robust Dimension Generalization with Entropy Normalization. While the distance features enable us to build a permutation-invariant attention module, they still rely on the assumption that the dimensions of feature- and label-spaces do not change across graphs. In practice, the dimensions of label spaces often vary a lot across datasets (e.g. 7 classes for Cora and 70 classes for FullCora). Consequently, the distance features have substantially varying scales (e.g. 7 versus 70) for different datasets, and the maximum and minimum distance become indistinguishable when the dimensions increase [5]. Both seriously hamper the generalization ability of the attention module trained on a single graph. While it is possible to offset the scale with a temperature hyperparameter specified for each dataset, this is contrary to our goal of graph foundation models. Instead, we would like a solution that automatically normalizes the distance features for any label dimension.
To this end, we employ entropy normalization, a technique widely used by manifold learning methods [17, 41] to adaptively determine the similarity between vectors. For node $u$, the asymmetric similarity feature between LinearGNN prediction $i$ and $j$ is given by:
@@
where $\sigma_{u}^{(i)}$ is the standard deviation of an isotropic multivariate Gaussian distribution, determined by matching the entropy of $p_{u}(j|i)$ with a hyperparameter $H$. Since the similarity features are functions of the distance features, they are also permutation invariant to the feature and label dimensions of the graph. Intuitively, this imposes a soft constraint on the number of LinearGNN predictions that are regarded as similar to the current prediction ${\bm{y}}_{u}^{(i)}$, which significantly reduces the gap between training and test features. We will verify the effectiveness of entropy normalization in Section 4.4.
3.3 Analysis of Time Complexity
One advantage of GraphAny is that it is more efficient than conventional graph neural networks (e.g. GCN [21], which is due to two reasons. First, LinearGNN leverages non-parameteric graph convolution operations, which can be preprocessed and cached for all nodes on a graph, reducing its inference time complexity to $O(|{\mathcal{V}}|)$. Second, the analytical solution used by LinearGNN eliminates all gradient steps, leading to a very low optimization cost for any graph $O(|{\mathcal{V}}_{L}|)$.
Table 1 shows the time complexity and total wall time of GCN, LinearGNN and GraphAny. The total wall time considers all training and inference time on 31 graphs. Even without any speed optimization in our implementation, GraphAny is 2.95$\times$ faster than GCN in total time. We believe the speedup can be even larger with dedicate implementations of GraphAny.
Table 1: Comparison of time complexity and wall time. $P$ is the number of epochs. Note that GCN has to be trained individually on each of the 31 graphs while GraphAny only needs 1 training graph.
@@ 3.3 Analysis of Time Complexity
GCN 0 $O(|{\mathcal{E}}|P)$ $O(|{\mathcal{E}}|)$ 18.80 min
LinearGNN $O(|{\mathcal{E}}|)$ $O(|{\mathcal{V}}_{L}|)$ $O(|{\mathcal{V}}|)$ 1.25 min (15.04$\times$)
GraphAny $O(|{\mathcal{E}}|)$ $O(|{\mathcal{V}}_{L}|)$ $O(|{\mathcal{V}}|)$ 6.37 min (2.95$\times$)
@@ 4 Experiments
In this section, we evaluate the performance of GraphAny against both transductive and inductive methods on 31 node classification datasets (Section B). We visualize the attention of GraphAny on different datasets, shedding light on what inductive knowledge our model has learned (Section 4.3). To provide a comprehensive understanding of GraphAny, we further conduct ablation studies on the proposed inductive attention and entropy normalization (Section 4.4,4.5).
@@ 4.1 Experimental Setup
Datasets. We have compiled a diverse collection of 31 node classification datasets from three sources: PyG [12], DGL [43], and OGB [19]. These datasets encompass a wide range of graph types including academic collaboration networks, social networks, e-commerce networks and knowledge graphs, with sizes varying from a few hundreds to a few millions of nodes. The number of classes across these datasets ranges from 2 to 70. Detailed statistics for each dataset are provided in Appendix B.
Implementation Details. For GraphAny, we employ 5 LinearGNNs with different graph convolution operations: ${\bm{F}}={\bm{X}}$ (Linear), ${\bm{F}}={\bm{A}}{\bm{X}}$ (LinearSGC1), ${\bm{F}}={\bm{A}}^{2}{\bm{X}}$ (LinearSGC2), ${\bm{F}}=({\bm{I}}-{\bm{A}}){\bm{X}}$ (LinearHGC1) and ${\bm{F}}=({\bm{I}}-{\bm{A}})^{2}{\bm{X}}$ (LinearHGC2), which cover identical, low-pass and high-pass spectral filters. In our experiments, we consider 4 GraphAny models trained separately on 4 datasets respectively: Cora (homophilic, small), Wisconsin (heterophilic, small), Arxiv (homophilic, medium), and Products (homophilic, large). The remaining 27 datasets are held out from these training sets, ensuring that the evaluations on these datasets are conducted in a fully inductive manner. More implementation details can be found at Appendix C. Code and data is available at this link222github.com/DeepGraphLearning/GraphAny.
Baselines. We compare GraphAny against conventional transductive methods, including MLP, GCN [21] and GAT [42]. These transductive models are trained separately for each dataset and can be considered as strong baselines for evaluating inductive models. Since there have been no inductive models for node classification on any graph, we consider non-parametric methods as baselines, including label propagation (LabelProp) [51] and all 5 LinearGNNs used in GraphAny. Note that these methods perform inductive inference but do not transfer any knowledge across graphs.
4.2 Inductive Inference on Any Graph
@@
Table 2 and Figure 5 show the results of GraphAny and baselines on 31 node classification datasets. We can see that our proposed LinearGNNs, with proper graph convolution operations, yield decent performance despite being non-parameteric. Specifically, LinearSGC2 is only 2.1% lower than GCN in accuracy, which coincides with the observation that SGC has performance similar to that of GCN [44]. However, LinearSGC2 uses an analytical solution for inference, which is about 15$\times$ faster than training a GCN from scratch on each dataset (Table  1).
By combining multiple LinearGNNs with learned attention, GraphAny further improves over LinearSGC2 by 3.0% accuracy, outperforming all transductive models. This improvement mainly comes from inductive generalization, since GraphAny achieves the best performance on the 27 held-out datasets rather than the 4 datasets we use to train GraphAny. An intriguing finding is the lack of significant difference between GraphAny trained on small datasets (e.g. Cora and Wisconsin) and large datasets (e.g. Arxiv and Products). We conjecture the reason is that even small datasets contain sufficiently diverse node patterns, allowing GraphAny to learn an effective attention mechanism.
4.3 Visualization of Learned Attention
@@
Category Method Cora Wisconsin Arxiv Products Held Out Avg. Total Avg. (27 graphs) (31 graphs) MLP 48.42$\pm$0.63 79.21$\pm$1.07 55.50$\pm$0.23 61.06$\pm$0.08 57.09 57.20 Transductive GCN 81.40$\pm$0.70 54.51$\pm$0.88 71.74$\pm$0.29 75.79$\pm$0.12 65.55 66.55 GAT 81.70$\pm$1.43 52.94$\pm$3.10 73.65$\pm$0.11 79.45$\pm$0.59 65.31 67.03 LabelProp 60.30$\pm$0.00 16.08$\pm$2.15 0.98$\pm$0.00 74.50$\pm$0.00 50.73$\pm$0.31 49.01$\pm$0.27 Linear 52.80$\pm$0.00 80.00$\pm$2.15 46.79$\pm$0.00 42.10$\pm$0.00 57.91$\pm$0.43 57.59$\pm$0.42 Non-parameteric LinearSGC1 74.30$\pm$0.00 45.49$\pm$13.96 55.33$\pm$0.00 56.58$\pm$0.00 62.69$\pm$0.24 62.08$\pm$0.48 LinearSGC2 78.20$\pm$0.00 57.64$\pm$1.07 59.58$\pm$0.00 62.92$\pm$0.00 64.38$\pm$0.48 64.41$\pm$0.39 LinearHGC1 22.50$\pm$0.00 64.32$\pm$2.15 22.92$\pm$0.00 15.00$\pm$0.00 37.01$\pm$0.20 36.26$\pm$0.23 LinearHGC2 23.80$\pm$0.00 56.08$\pm$4.29 20.65$\pm$0.00 13.39$\pm$0.00 35.62$\pm$0.68 34.70$\pm$0.55 GraphAny (Cora) 80.18 $\pm$0.13 61.18$\pm$5.08 58.62$\pm$0.05 61.60$\pm$0.10 67.24$\pm$0.23 67.00$\pm$0.14 Inductive GraphAny (Wisconsin) 77.82$\pm$1.15 71.77$\pm$5.98 57.79$\pm$0.56 60.28$\pm$0.80 67.31$\pm$0.38 67.26$\pm$0.20 (training set) GraphAny (Arxiv) 79.38$\pm$0.16 65.10$\pm$3.22 58.68$\pm$0.17 61.31$\pm$0.20 67.65$\pm$0.31 67.46$\pm$0.27 GraphAny (Products) 79.36$\pm$0.23 65.89$\pm$2.23 58.58$\pm$0.11 61.19$\pm$0.23 67.66$\pm$0.39 67.48$\pm$0.33
@@
Figure 6: Normalized performance of LinearGNNs (left, best as 1) and attention weights of GraphAny trained on Wisconsin (middle) and Arxiv (right) respectively. The best and second best LinearGNN performance and attention weights for each dataset are highlighted with red and purple rectangles respectively. Our inductive attention can identify the best-performing LinearGNN for most datasets.
@@
To understand how LinearGNNs are combined in GraphAny by the inductive attention, we visualize the attention weights of GraphAny (Wisconsin) and GraphAny (Arxiv) on all datasets, averaged across nodes. For reference, we also visualize the performance of each individual LinearGNN on all datasets. As shown in Figure 6, we can see that LinearSGC2 is optimal for half of the datasets, while the other half prefers LinearHGC1, Linear or LinearSGC1. In most cases, GraphAny successfully identifies the optimal LinearGNN within its top-2 attention weights, with Hits@2 being 0.65 and 0.77 for GraphAny trained on Wisconsin and Arxiv respectively.
@@
Interestingly, there is a distinction between the attention distributions learned on the two datasets: Wisconsin leads a relatively balanced distribution of attention across 5 LinearGNNs, while Arxiv prefers a more focused distribution of attention, favoring low-pass filters like LinearSGC1 and LinearSGC2. This reflects the nature of these training sets: Wisconsin is a heterophilic dataset where all LinearGNNs are reasonably good, but Arxiv is a homophilic dataset where Linear, LinearSGC1 and LinearSGC2 have better performance (Table 2). Note that this does not necessarily mean GraphAny cannot attend to a LinearGNN that is not preferred by the training graph at node level, but the averaged attention at graph level does correlate with the distribution of the training graph.
4.4 Ablation Study on Entropy Normalization
One key component in GraphAny is the entropy normalization technique applied to distance features. To provide insights on how entropy normalization improves the dimension generalization ability of our method, we plot the feature distribution between the predictions of LinearSGC2 and other LinearGNNs on Cora and FullCora, which have similar semantics but different number of classes. As shown in Figure 7, there is a large difference in the scale of unnormalized feature distributions. By constrast, with entropy normalization, the feature distributions become very similar on these datasets, which facilitates inductive generalization across different datasets.

### v3

@@ 3.2 Learning Inductive Attention over LinearGNN Predictions
While there are various ways to parameterize this attention module, finding an inductive solution is non-trivial since neural networks can easily fit the information specific to the training graph. We notice a necessary property for fully-inductive functions is that it should be robust to transformations on features and labels, such as permutation or masking on some dimensions (Figure 4). This requires our attention module, to be permutation invariant and robust to dimension changes, which motivates our design of distance features and entropy normalization respectively.
Permutation-Invariant Attention with Distance Features.
We would like to design an attention module that is permutation invariant (Bronstein et al., 2021) along the data dimension. [^1] It is important to distinguish between domain symmetry (in this case, permutation of the nodes of the graph) and data symmetry (permutation of the feature and label dimensions). Invariance to domain symmetry (node permutations) is provided by design in our LinearGNN. Consider a new graph generated by permuting feature and label dimensions of the training graph. We expect our attention output to be invariant to these permutations in order to generate the same prediction as for the unpermutated training set.
@@
which is invariant to both feature and label-permutation matrices ${\bm{P}}$ and ${\bm{Q}}$. Generally, any feature that is a linear combination of dot products between LinearGNN predictions is also permutation invariant, e.g. Euclidean distance or Jensen-Shannon divergence. For a set of $t$ LinearGNN predictions $\hat{{\bm{y}}}_{u}^{(1)},\hat{{\bm{y}}}_{u}^{(2)},\cdots,\hat{{\bm{y}}}_{u}^{(t)}$ on a single node $u$, we construct the following $t(t-1)$ permutation-invariant features to capture their squared distances:
@@
Robust Dimension Generalization with Entropy Normalization.
Figure 5: Comparison of Euclidean distances (the first row) and entropy-normalized (the second row) features between five channels: ${\bm{F}}={\bm{X}}$ (Linear), ${\bm{F}}=\bar{{\bm{A}}}{\bm{X}}$ (LinearSGC1), ${\bm{F}}=\bar{{\bm{A}}}^{2}{\bm{X}}$ (LinearSGC2), ${\bm{F}}=({\bm{I}}-\bar{{\bm{A}}}){\bm{X}}$ (LinearHGC1) and ${\bm{F}}=({\bm{I}}-\bar{{\bm{A}}})^{2}{\bm{X}}$ (LinearHGC2) with $\bar{{\bm{A}}}$ denoting the row normalized adjaceny matrix. Entropy-normalized features are on the same scale and exhibit transferrable patterns across datasets.
While the distance features for inductive attention ensure a permutation-invariant attention module, the distance features are known to suffer from the curse of dimensionality, where distances between vectors with larger dimensions have smaller scales (Beyer et al., 1999). This will hamper the generalization performance when the dimensions of label spaces vary across training and inference graphs (e.g. training on 7 classes for Cora and inference on 70 classes for FullCora). Empirically, as shown in the first row of Figure 5, the scale of Euclidean distance distributions decreases drastically when the number of classes increases, hampering the generalization ability of the attention module trained on a single graph.
A naive approach is to normalize the distance-feature distributions using a hyperparameter (e.g., temperature). However, due to the varying neighbor patterns across graphs (or even nodes (Luan et al., 2022)), a single hyperparameter may not be suitable for all nodes and graphs. Instead, we propose an adaptive solution that normalizes distance features to a consistent scale. To achieve this, we employ entropy normalization, a technique commonly used in manifold learning (Hinton and Roweis, 2002, van der Maaten and Hinton, 2008) to adaptively determine the similarity features. For node $u$, the asymmetric similarity feature between LinearGNN predictions $i$ and $j$ is defined as:
@@
where $\sigma_{u}^{(i)}$ is the standard deviation of an isotropic multivariate Gaussian, determined by matching the entropy of distance distributions $P_{u}^{(i)}=\{p_{u}(j\mid i)\mid j\in[1,t]\}$ to a fixed hyperparameter $H$. Since the similarity features are derived from distance features, they are also permutation-invariant to the feature and label dimensions of the graph. Intuitively, this imposes a soft constraint on the number of LinearGNN predictions considered similar to $\hat{{\bm{y}}}_{u}^{(i)}$, significantly reducing the gap between training and test features. As shown in the second row of Figure 5, entropy-normalized feature distributions are on consistent scales across datasets. Additionally, we observe that different types of homophilic graphs (e.g., the citation graph Cora and the e-commerce graph Product) exhibit similar entropy-normed features. We will further verify the effectiveness of entropy normalization empirically in Section 4.4.
3.3 Efficient Training and Fully-inductive Inference
In this section, we summarize how GraphAny utilizes the techniques introduced in the previous section and derive an efficient fully-inductive node classification model. As shown in Figure 3, given any graph, GraphAny first utilizes $t$ LinearGNNs to provide basic predictions with different channels. Then, entropy-normalized features are computed based on the distances between these predictions, leading to $t(t-1)$-dimensional features. Further, an inductive attention module $f_{\theta}$ (e.g. MLP) is used to compute the attention scores for fusing different predictions into a final prediction (Eq. 5). Since the only trainable module of GraphAny lies in the inductive attention module $f_{\theta}:\mathbb{R}^{t(t-1)}\rightarrow\mathbb{R}^{t}$, which is independent to feature dimension $d$ and label dimension $c$, GraphAny enjoys fully-inductive inference on any graph with arbitrary feature and label dimensions.
One advantage of GraphAny is that it is more efficient than conventional graph neural networks (e.g. GCN (Kipf and Welling, 2017), which is due to two reasons. First, LinearGNN leverages non-parameteric graph convolution operations, which can be preprocessed and cached for all nodes on a graph, reducing the optimization complexity to $O(|{\mathcal{V}}_{L}|)$ compared with the complexity of $O(|\mathcal{E}|)$ for standard GNNs. Second, once trained, GraphAny is ready to generalize to arbitrary graphs, eliminating the need for gradient descent on test graphs.
Table 1 shows the time complexity and total wall time of GCN, LinearGNN and GraphAny. The total wall time considers all training and inference time on 31 graphs. Even without any speed optimization in our implementation, GraphAny is 2.95$\times$ faster than the optimized DGL’s (Wang et al., 2020) GCN implementation in total time. We believe the speedup can be even larger with dedicate implementations of GraphAny.
Table 1: Comparison of time complexity and wall time. Note that GCN has to be trained individually on each of the 31 graphs while GraphAny only needs 1 training graph.
@@ 3.3 Efficient Training and Fully-inductive Inference
GCN 0 $O(|{\mathcal{E}}|)$ $O(|{\mathcal{E}}|)$ 18.80 min
LinearGNN $O(|{\mathcal{E}}|)$ $O(|{\mathcal{V}}_{L}|)$ $O(|{\mathcal{V}}_{U}|)$ 1.25 min (15.04$\times$)
GraphAny $O(|{\mathcal{E}}|)$ $O(|{\mathcal{V}}_{L}|)$ $O(|{\mathcal{V}}_{U}|)$ 6.37 min (2.95$\times$)
@@ 4 Experiments
In this section, we evaluate the performance of GraphAny against both transductive and inductive methods on 31 node classification datasets (details in Appendix B). We visualize the attention of GraphAny on different datasets, shedding light on what inductive knowledge our model has learned (Section 4.3). To provide a comprehensive understanding of the proposed techniques of GraphAny, we further conduct ablation studies on the entropy-normalized feature and attention parameterization in Section 4.4. More training and implementation details are provided in Appendix C.
@@ 4.1 Experimental Setup
Datasets. We have compiled a diverse collection of 31 node classification datasets from three sources: PyG (Fey and Lenssen, 2019), DGL (Wang et al., 2020), and OGB (Hu et al., 2021b). These datasets encompass a wide range of graph types including academic collaboration networks, social networks, e-commerce networks and knowledge graphs, with sizes varying from a few hundreds to a few millions of nodes. The number of classes across these datasets ranges from 2 to 70. Detailed statistics for each dataset are provided in Appendix B.
Implementation Details. For GraphAny, we employ 5 LinearGNNs with different graph convolution operations: ${\bm{F}}={\bm{X}}$ (Linear), ${\bm{F}}=\bar{{\bm{A}}}{\bm{X}}$ (LinearSGC1), ${\bm{F}}=\bar{{\bm{A}}}^{2}{\bm{X}}$ (LinearSGC2), ${\bm{F}}=({\bm{I}}-\bar{{\bm{A}}}){\bm{X}}$ (LinearHGC1) and ${\bm{F}}=({\bm{I}}-\bar{{\bm{A}}})^{2}{\bm{X}}$ (LinearHGC2) with $\bar{{\bm{A}}}$ denoting the row normalized adjaceny matrix, which cover identical, low-pass and high-pass spectral filters. In our experiments, we consider 4 GraphAny models trained separately on 4 datasets respectively: Cora (homophilic, small), Wisconsin (heterophilic, small), Arxiv (homophilic, medium), and Products (homophilic, large). The remaining 27 datasets are held out from these training sets, ensuring that the evaluations on these datasets are conducted in a fully-inductive manner. More implementation details can be found at Appendix C.
Baselines. As there are no existing fully-inductive node classification baselines, we include non-parametric methods (models without learnable parameters) like label propagation (Zhu and Ghahramani, 2002) and the five LinearGNNs used in GraphAny. While these methods perform inductive inference, they do not transfer knowledge across graphs. Additionally, we compare GraphAny with transductive models, including MLP, GCN (Kipf and Welling, 2017), and GAT (Velickovic et al., 2018). These models are trained separately on each dataset and serve as strong baselines for inductive models, as they benefit from backpropagation on labeled nodes and hyperparameter tuning on validation sets of the test dataset.
4.2 Performance of Inductive Node Classification
@@
Table 2 presents the results of GraphAny and various baselines on 31 node classification datasets (complete results for each dataset are provided in Appendix D). Our proposed LinearGNNs, despite being non-parametric, demonstrate competitive performance. Notably, LinearSGC2, a linear model with a two-hop graph convolution layer, achieves only 2.1% lower accuracy than GCN, which aligns with previous findings that SGC performs comparably to GCN (Wu et al., 2019). Moreover, LinearSGC2 leverages an analytical solution for inference, making it approximately 15$\times$ faster than training a GCN from scratch on each dataset (see Table 1). Additionally, we observe that the optimal LinearGNN model differs across datasets, highlighting that no single graph convolution kernel is universally effective for all graphs.
As for GraphAny, which is trained on just 1 of the 31 graphs, it significantly outperforms LinearGNNs and even slightly surpasses transductive baselines that are individually trained on all 31 graphs. This improvement is primarily driven by inductive generalization, as GraphAny achieves its strongest performance on the 27 held-out (fully-inductive) datasets rather than the 4 training (transductive) datasets. A closer examination of Figure 6 shows that GraphAny performs well on both homophilic and heterophilic graphs in an inductive manner. We attribute this to the inductive attention module, which adaptively fuses predictions from different graph convolution kernels for each node. Interestingly, we also observe minimal performance differences between GraphAny when trained on small datasets (e.g., Cora and Wisconsin) and large datasets (e.g., Arxiv and Products). We hypothesize that even small datasets contain sufficiently diverse local node patterns (e.g., homophily and heterophily) (Luan et al., 2022), enabling GraphAny to learn an effective node-level attention mechanism that generalizes well with a limited number of nodes (e.g., 120). A detailed analysis of the attention module is provided in Section 4.3.
4.3 Visualization of the Inductive Attention
@@
Category Method Cora Wisconsin Arxiv Products Held Out Avg. Total Avg. (27 graphs) (31 graphs) MLP 48.42$\pm$0.63 66.67$\pm$3.51 55.50$\pm$0.23 61.06$\pm$0.08 57.09 57.20 Transductive GCN 81.40$\pm$0.70 37.25$\pm$1.64 71.74$\pm$0.29 75.79$\pm$0.12 65.55 66.55 GAT 81.70$\pm$1.43 52.94$\pm$3.10 73.65$\pm$0.11 79.45$\pm$0.59 65.31 67.03 LabelProp 60.30$\pm$0.00 16.08$\pm$2.15 0.98$\pm$0.00 74.50$\pm$0.00 50.73$\pm$0.31 49.01$\pm$0.27 Linear 52.80$\pm$0.00 80.00$\pm$2.15 46.79$\pm$0.00 42.10$\pm$0.00 57.91$\pm$0.43 57.59$\pm$0.42 Non-parameteric LinearSGC1 74.30$\pm$0.00 45.49$\pm$13.96 55.33$\pm$0.00 56.58$\pm$0.00 62.69$\pm$0.24 62.08$\pm$0.48 LinearSGC2 78.20$\pm$0.00 57.64$\pm$1.07 59.58$\pm$0.00 62.92$\pm$0.00 64.38$\pm$0.48 64.41$\pm$0.39 LinearHGC1 22.50$\pm$0.00 64.32$\pm$2.15 22.92$\pm$0.00 15.00$\pm$0.00 37.01$\pm$0.20 36.26$\pm$0.23 LinearHGC2 23.80$\pm$0.00 56.08$\pm$4.29 20.65$\pm$0.00 13.39$\pm$0.00 35.62$\pm$0.68 34.70$\pm$0.55 GraphAny (Cora) 80.18 $\pm$0.13 61.18$\pm$5.08 58.62$\pm$0.05 61.60$\pm$0.10 67.24$\pm$0.23 67.00$\pm$0.14 Inductive GraphAny (Wisconsin) 77.82$\pm$1.15 71.77$\pm$5.98 57.79$\pm$0.56 60.28$\pm$0.80 67.31$\pm$0.38 67.26$\pm$0.20 (training set) GraphAny (Arxiv) 79.38$\pm$0.16 65.10$\pm$3.22 58.68$\pm$0.17 61.31$\pm$0.20 67.65$\pm$0.31 67.46$\pm$0.27 GraphAny (Products) 79.36$\pm$0.23 65.89$\pm$2.23 58.58$\pm$0.11 61.19$\pm$0.23 67.66$\pm$0.39 67.48$\pm$0.33
@@
Figure 7: Normalized performance of LinearGNNs (left; best as 1.00) and attention weights of GraphAny trained on Wisconsin (middle) and Arxiv (right) respectively. The best and second best LinearGNN performance and attention weights for each dataset are highlighted with red and purple rectangles respectively. The learned inductive attention of GraphAny successfully identifies the best-performing LinearGNN for most datasets.
@@
To understand how LinearGNNs are combined in GraphAny by the inductive attention, we visualize the attention weights of GraphAny (Wisconsin) and GraphAny (Arxiv) on all datasets, averaged across nodes. For reference, we also visualize the performance of each individual LinearGNN on all datasets. As shown in Figure 7, we can see that half of the datasets are homophilic with LinearSGC2 being the optimal LinearGNN, while the other half prefers LinearHGC1, Linear or LinearSGC1. In most cases, GraphAny successfully identifies the optimal LinearGNN within its top-2 attention weights, with Hits@2 being 0.65 and 0.77 for GraphAny trained on Wisconsin and Arxiv respectively.
@@
We hypothesis that this amazing inductive performance comes from the inductive entropy-normed distance feature we derived, where homophilic and heterophilic graphs share different patterns (see the second row in Figure 5). Interestingly, there is a distinction between the attention distributions when training on different datasets: GraphAny-Wisconsin leads a relatively balanced distribution of attention across 5 LinearGNNs, while Graph-Arxiv prefers a more focused distribution of attention, favoring low-pass filters like LinearSGC1 and LinearSGC2. This reflects the nature of these training sets: As shown in the left part of Figure 7, all LinearGNN channels in Wisconsin are reasonably good, indicating diverse message-passing pattern exists, but Arxiv is a homophilic dataset where Linear, LinearSGC1 and LinearSGC2 have better performance (Table 2). These different message-passing patterns are learned and used by GraphAny to generate inductive attention scores.
4.4 Ablation Studies